### v3

# Enhancing Agentic Textual Graph Retrieval with Synthetic Stepwise Supervision

Abstract
Integrating textual graphs into Large Language Models (LLMs) is promising for complex graph-based QA. However, a key bottleneck is retrieving informative yet compact subgraphs that fit the LLM context. Existing retrievers often struggle, relying either on shallow embedding similarity or costly interactive policies that require excessive supervision. To address these challenges, we introduce Graph-S3, an agentic textual graph reasoning framework featuring an LLM-based retriever trained with synthetic stepwise supervision. Rather than relying on final answer rewards—which often yield sparse and unstable signals—we optimize the retriever by evaluating each step against offline-extracted golden subgraphs. Our approach distills golden subgraphs via a specialized data synthesis pipeline to formulate dense rewards, facilitating a two-stage training scheme that effectively learns the interactive graph exploration policy. Based on extensive experiments on three common datasets in comparison with several strong baselines, our approach achieves an average improvement of 15.6% in accuracy and 17.2% in F1 score. The advantage is even higher in more complicated multi-hop reasoning tasks.
Enhancing Agentic Textual Graph Retrieval with Synthetic Stepwise Supervision
[…]
11footnotetext: Equal contribution.22footnotetext: Corresponding author.33footnotetext: Work done during internship at AIR, Tsinghua University.
1 Introduction
Textual graphs are widely employed for structured knowledge representation in domains like question answering and scientific discovery (Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2025a). By explicitly modeling multi-hop relations, they enable interpretable, compositional reasoning that is difficult to achieve with unstructured text (Chen et al., 2020; Hogan et al., 2021; Zou, 2020). While early methods were constrained by inflexible symbolic inference and high annotation costs (Yih et al., 2016), the semantic capabilities of Large Language Models (LLMs) have alleviated these limitations (Chang et al., 2024; Su et al., 2025). Consequently, integrating LLMs with textual graphs has become a prominent paradigm for general-purpose graph understanding and QA (Lewis et al., 2020; Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2025a; Jin et al., 2024; Chai et al., 2023).
Figure 1: An illustration of agentic textual graph retrieval and question answering.
LLM agents accessing graphs via tool calls surpass simple retrieval (Jiang et al., 2024; Yang et al., 2024; Ji et al., 2024; Xu et al., 2026), as illustrated in Figure 1 where an agent iteratively collects information. However, training such agents via supervised fine-tuning (SFT) often collapses the action space, favoring memorization over generalizable policies (Chu et al., 2025; Li et al., 2024). While outcome-supervised reinforcement learning (RL) offers a potential solution by exploring reasoning trajectories (Lightman et al., 2023; Paolo et al., 2024), it faces significant hurdles in this domain: the action space in real-world textual graphs is often too vast for efficient exploration, and outcome-based rewards are sparse and noisy. For instance, redundant or erroneous retrieval steps may coincidentally lead to the correct final answer, rendering outcome supervision unreliable for assessing reasoning quality (Liu et al., 2023; Rengarajan et al., 2022).
To overcome this limitation, we introduce a synthetic stepwise supervision scheme that provides explicit feedback at every decision step, ensuring that the model is guided not only by the correctness of the final answer but also by the quality of intermediate actions. The key idea is to guide each graph retrieval step with golden subgraphs offline extracted from the target graph. Specifically, we propose an automated pipeline to construct the golden subgraphs for reward computation. We first generate a large amount of subgraph candidates through random and LLM-guided exploration, and filter the candidates based on information sufficiency, i.e. whether they are able to produce correct answers with LLMs. These successful exploratory trajectories are used for SFT, providing the retriever with basic navigation ability as a warm-up stage. Then we further refine the subgraphs to enhance information conciseness by iteratively pruning redundant content while preserving answer consistency. With these refined subgraphs, each online graph retrieval action can be associated with an explicit stepwise reward based on its contribution to the golden subgraphs. The combined two-step training pipeline guides the retriever to improve reasoning decisions over long action chains.
Extensive experiments have demonstrated the effectiveness of our synthetic stepwise supervision approach. For example, while retrieving only 11.44% of the triples, Graph-S3 achieves an average improvement of 15.6% in accuracy and 17.2% in F1 score across the WebQSP, CWQ, and MetaQA benchmarks.
In summary, the main contributions of this work are as follows:
(1) We propose an automatic pipeline for synthesizing high-quality stepwise supervision data for interactive graph retrieval, addressing the scarcity of fine-grained training signals in this field.
(2) We design a two-stage training paradigm tailored for graph reasoning: SFT on raw synthetic trajectories to bootstrap basic navigation ability, followed by RL with synthetic stepwise rewards on refined trajectories to provide explicit feedback and strengthen reasoning strategies.
(3) Experimental results demonstrate that Graph-S3 achieves state-of-the-art performance on the WebQSP, CWQ, and MetaQA datasets with accurate and compact graph retrieval.
Figure 2: An Overview of our data synthesis pipeline and training scheme.
2 Related Work
2.1 Graph Retrieval Methods
Graph retrieval approaches include similarity-based, GNN-based, and LLM-based methods (Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2025a; Zhu et al., 2025; Han et al., 2024; Zhou et al., 2026), but most perform one-shot retrieval and often return redundant or incomplete subgraphs. Recent interactive frameworks (Jiang et al., 2024; Ji et al., 2024; Yang et al., 2024) allow iterative exploration, yet their training predominantly relies on imitation of language patterns or outcome-based supervision, which provides only coarse feedback and limits stable multi-hop reasoning. In contrast, our work employs RL with synthetic stepwise rewards and a scalable data synthesis pipeline to provide supervision for interactive retrieval.
2.2 Stepwise RL for Graph Reasoning
Recent advances such as OpenAI o1 and DeepSeek-R1 (Jaech et al., 2024; Guo et al., 2025) demonstrate the effectiveness of RL in strengthening multi-step reasoning, enabling models to perform longer chains of thought with improved reliability in domains like mathematics and programming (Guo et al., 2024a; El-Kishky et al., 2025; Li et al., 2026; Xu et al., 2025). In contrast, applying RL to textual graphs remains limited, partly due to the lack of fine-grained supervision data (Zhang et al., 2025a; Yao et al., 2025; Liu et al., 2025). RL-based graph agents  (Das et al., 2017; Cui et al., 2025) relied on sparse outcome rewards, making credit assignment across reasoning trajectories difficult. Although recent efforts have introduced reasoning-structured datasets (Pahilajani et al., 2024), high-quality stepwise supervision for graph-based RL remains scarce and difficult to construct at scale. These observations highlight the need for scalable approaches that can provide fine-grained supervision signals and support stable optimization for interactive graph retrieval.
3 Method
We present our agentic retrieval framework, designed to equip large language models with robust graph reasoning capabilities through stepwise supervision, a two-stage training paradigm, and an interactive retrieval strategy. As illustrated in Figure 2, the framework comprises three main components. First, we construct an automatic data synthesis pipeline that leverages GPT-4o (OpenAI, 2024) to generate diverse exploratory trajectories, which are subsequently refined into high-quality stepwise supervision data. This addresses the scarcity of fine-grained training signals for graph-based RL. Second, we adopt a two-stage training paradigm: SFT on raw synthetic trajectories provides a warm-up initialization for basic graph navigation, while RL with synthetic stepwise rewards on refined trajectories supplies explicit feedback at each decision step, stabilizing optimization and strengthening reasoning strategies. Finally, during inference, the retriever operates under an interactive retrieval mechanism that conducts stepwise, structure-aware exploration of the textual graph, thereby reducing redundancy and mitigating incomplete retrieval.
3.1 Data Synthesis
Existing LLMs are not pretrained on graph-structured data (Zhang et al., 2025b), which significantly limits their performance on graph reasoning tasks. As a result, effective training requires high-quality supervision to cultivate graph comprehension and reasoning capabilities. However, constructing such datasets is notoriously expensive, since it often relies on manual annotation by domain experts (Choubey et al., 2024), leading to a persistent scarcity of high-quality graph reasoning data. To address this issue, we design a pipeline that automatically synthesizes graph reasoning trajectories. Specifically, we first define a set of actions that enable structured interaction with the graph. Then, we leverage GPT-4o to perform these actions and collect valid action–response pairs, which form exploratory trajectories. The raw trajectories are directly used for SFT to provide the model with basic navigation ability, while a refinement step prunes redundant detours and preserves all answer-consistent trajectories, producing high-quality stepwise supervision data for RL.
3.1.1 Action Space for Graph Exploration
Given a textual graph $\mathcal{G}=\{t_{i}\}_{i=1}^{m}$, where each triple $t_{i}=(e_{h}^{i},r^{i},e_{t}^{i})\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ consists of a head entity $e_{h}^{i}\in\mathcal{E}$, a relation $r^{i}\in\mathcal{R}$, and a tail entity $e_{t}^{i}\in\mathcal{E}$. Here, $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and relations, respectively. To enable stepwise exploration over $\mathcal{G}$, we define the retriever’s action space as consisting of three types of operations. For clarity, we use $(x,r,y)$ to denote a generic triple in $\mathcal{G}$, where $x,y\in\mathcal{E}$ and $r\in\mathcal{R}$.
Explore Entity: This operation expands the local neighborhood of a given entity by retrieving all directly connected triples in $\mathcal{G}$. Formally, for a target entity $x\in\mathcal{E}$, the operation is defined as
$\text{Explore}(x)=$ $\{(x,r,y)|(x,r,y)\in\mathcal{G}^{p}\}$
$\cup\{(y,r,x)|(y,r,x)\in\mathcal{G}^{p}\}$ (1)
where $r\in\mathcal{R}$ and $y\in\mathcal{E}$ denote relations and neighboring entities, respectively. The retrieved triples are added to the perception window $\mathcal{G}^{p}$ for subsequent reasoning steps.
Choose Relation: The perception window $\mathcal{G}^{p}$ obtained from the Explore action may still contain many irrelevant triples. To avoid introducing redundant context into the LLM, this operation prunes $\mathcal{G}^{p}$ into a query-relevant subgraph $\mathcal{G}^{sub}$:
$\begin{split}\operatorname{Choose}(q,\mathcal{G}^{p})=\bigl\{&(x,r,y)\in\mathcal{G}^{p}\big|\\ &F(q,(x,r,y))=1\bigr\}.\end{split}$ (2)
Here, $q$ is the query, $(x,r,y)$ denotes a triple in $\mathcal{G}^{p}$, and $F(q,(x,r,y))$ is the relevance function learned by the retriever to decide whether the triple should be preserved. $F$ outputs $1$ if the triple is judged relevant to the query and 0 otherwise.
Finish: This operation indicates that the retriever has gathered sufficient evidence in the current subgraph $\mathcal{G}^{sub}$ to answer the query $q$. Once invoked, the exploration process terminates, and $(q,\mathcal{G}^{sub})$ is used to produce the final answer:
$\text{Finish}(q,\mathcal{G}^{sub})=\text{Answer}(q,\mathcal{G}^{sub}),$ (3)
where $\text{Answer}(q,\mathcal{G}^{sub})$ denotes answering query $q$ based on the retrieved subgraph.
3.1.2 Graph Reasoning Data Synthesis
Given the defined action space, we synthesize reasoning trajectories by letting a behavior model (GPT-4o) interact with the graph. We formalize the generation process as a Markov decision process (MDP) with deterministic transitions defined by the graph action space. Each trajectory consists of multiple decision steps, and is later decomposed into step-level training instances for SFT and RL.
State. We define each state as a tuple of four components: $s_{t}=\big(q,\;\mathcal{G}^{p}_{t},\;\mathcal{G}^{sub}_{t},\;h_{t}\big)$, where $q$ is the query, $\mathcal{G}^{p}_{t}$ the perception window aggregated by Explore, and $\mathcal{G}^{sub}_{t}$ the focused subgraph maintained by Choose, and $h_{t}=(a_{1},\dots,a_{t-1})$ the action history up to step $t$. Including $h_{t}$ allows the retriever to condition its decisions not only on the current graph view but also on the reasoning trajectory already taken.
Action. The parameterized action space is $\mathcal{A}$, defined in 3.1.1. During the synthesis phase, the behavior model selects $a_{t}\in\mathcal{A}$ given $s_{t}$, and additionally produces a natural language reasoning trace that explains the choice of action.
Transition. The execution of action $a_{t}$ triggers a state transition $s_{t}\to s_{t+1}$ by updating the action history $h_{t+1}\leftarrow h_{t}\cup\{a_{t}\}$ and modifying the graph components based on the action type. Specifically, if $a_{t}=\textsc{Explore}(x_{t})$, the agent expands the partial graph $\mathcal{G}^{p}_{t}$ by merging it with the newly explored neighbors. If $a_{t}=\textsc{Choose}$, the agent updates the subgraph $\mathcal{G}^{sub}_{t}$ by filtering relevant triples from $\mathcal{G}^{p}_{t}$ via the selection function. The episode terminates when $a_{t}=\textsc{Finish}$ or the time step $t$ reaches the maximum limit $T_{\max}$.
Answer and retention. Upon termination at step $T$, we produce an answer $\hat{y}\;=\;\text{Answer}\!\big(q,\mathcal{G}^{sub}_{T}\big)$ and keep the trajectory $\tau\;=\;\big\{(s_{t},a_{t})\big\}_{t=1}^{T}$ only if $\hat{y}$ matches the set of ground-truth answers. After that, we use the raw action labels for SFT dataset: $\mathcal{D}_{\text{SFT}}\;=\;\big\{(s_{t},a_{t})\big\}_{\tau,\,t}$. In the next subsection, we further refine trajectories to obtain stepwise supervision signals for RL.
3.1.3 Trajectory Refinement for RL
While SFT directly benefits from raw trajectories, RL requires more concise training signals (Yue et al., 2025). Filtering trajectories only by final answer correctness often produces redundant exploration steps, introducing noise and inefficiency during policy optimization. To address this, we introduce a refinement procedure that removes unnecessary detours while preserving all answer-consistent trajectories, thereby yielding shorter and cleaner trajectories for RL.
Let the set of retained raw trajectories be $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$, where each $\tau_{i}=\{(s_{t},a_{t})\}_{t=1}^{T_{i}}$ is a sequence of state–action pairs terminating at step $T_{i}$. We define a refinement operator $\mathcal{R}$ that maps the raw set $\mathcal{T}$ into a refined set $\mathcal{T}^{*}$:
$\mathcal{T}^{*}=\mathcal{R}(\mathcal{T}).$ (4)
For each $\tau_{i}$, the refinement identifies the shortest feasible subsequence $\tau_{i}^{*}$ that still leads to the same correct final answer:
$\tau_{i}^{*}=\arg\min_{\tau\in\mathcal{F}_{i}}|\tau|,$ (5)
where $\mathcal{F}_{i}$ is the set of all feasible answer-consistent trajectories equivalent to $\tau_{i}$. Let $s_{T}(\tau)$ denotes the final state and $y(\tau)$ denotes the answer.
$\mathcal{F}_{i}=\big\{\tau\mid s_{T}(\tau)=s_{T}(\tau_{i})\land y(\tau)=y(\tau_{i})\big\}.$ (6)
Thus, the refined dataset $\mathcal{T}^{*}$ retains trajectories that are semantically equivalent to the originals but stripped of redundant exploration steps. Each refined trajectory $\tau_{i}^{*}$ is then decomposed into step-level supervision signals by attaching a rule-based stepwise reward $\ell_{t}\in[0,1]$ to each action $a_{t}$, indicating its correctness within the reasoning trajectory. Formally, the RL training dataset is constructed as $\mathcal{D}_{\text{RL}}=\big\{(s_{t},a_{t},\ell_{t})\big\}_{\tau^{*},\,t}$. This ensures that RL receives concise stepwise supervision signals, improving both stability and efficiency of training.
3.2 Training Stage
To enhance the model’s graph comprehension and reasoning capabilities, we adopt a two-stage fine-tuning approach. The first stage uses SFT with synthesized data to establish foundational abilities. The second stage employs GRPO with trajectory refinement, leveraging RL’s proven effectiveness in enhancing reasoning and exploration efficiency (Yue et al., 2025).
3.2.1 Stage I: Supervised fine-tuning
For each step $t$, let $s_{t}=(q,\mathcal{G}^{p}_{t},\mathcal{G}^{sub}_{t},h_{t})$ be the serialized state, and let $y_{t}=(y_{t}^{1},\dots,y_{t}^{L_{t}})$ be the target token sequence that concatenates the natural-language thought process and the action specification. Denote by $\mathcal{I}(s_{t})$ the textual serialization of the state. The training loss of SFT is defined as
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(s_{t},y_{t})\\ \sim\mathcal{D}_{\text{SFT}}\end{subarray}}\Bigg[\sum_{l=1}^{L_{t}}\log\pi_{\theta}\big(y_{t}^{\,l}\mid\mathcal{I}(s_{t}),\,y_{t}^{<l}\big)\Bigg].$ (7)
where $\pi_{\theta}(y_{t}^{\,l}\mid\mathcal{I}(s_{t}),y_{t}^{<l})$ denotes the probability assigned by the model to the $l$-th token given the serialized state and the previously generated tokens.
3.2.2 Stage II: RL with stepwise rewards
For reward design, existing approaches predominantly rely on outcome-based reward signals, which have demonstrated remarkable effectiveness in domains such as mathematical reasoning and code generation. However, prior studies (Wang et al., 2025; Choudhury, 2025; Deng et al., 2024) have shown that in relatively complex scenarios such as graph retrieval, conventional outcome-based reward signals tend to be overly sparse. This sparsity hampers effective credit assignment to early-stage actions, ultimately resulting in inefficient learning over long action chains. This observation motivates our adoption of process-level rewards in the training process, where the reward signal is determined by the contribution of each current action to the golden subgraphs.
Specifically, each step $t$ is associated with a process-level rule-based reward $\ell_{t}$ that provides graded feedback according to the quality of the predicted action:
$\ell_{t}=\begin{cases}0,&\text{if }\mathrm{Inv}(a_{t}),\\ c_{1},&\text{if }\mathrm{Fmt}(a_{t})=1\land\mathrm{Corr}(a_{t})=0,\\ c_{2},&\text{if }\mathrm{Part}(a_{t})=1,\\ 1.0,&\text{if }a_{t}=a_{t}^{*}.\end{cases}$ (8)
Here, $\mathrm{Inv}(\cdot)$, $\mathrm{Fmt}(\cdot)$, $\mathrm{Corr}(\cdot)$, and $\mathrm{Part}(\cdot)$ denote the deterministic functions for checking validity, format correctness, action correctness, and partial correctness, respectively. $c_{1}$ and $c_{2}$ are dataset-specific hyperparameters. Consequently, we adopt the reward function shown in Eq. 8 to train our model using the GRPO method.
3.3 Interactive Retriever
At inference time, Graph-S3 interacts with the textual graph through the defined action space. Unlike single-pass retrieval methods that return large subgraphs, our approach performs stepwise exploration by balancing EXPLORE and CHOOSE actions, terminating with FINISH when sufficient evidence is gathered. This interactive process enables precise control over retrieval depth while minimizing redundancy, producing concise subgraphs for reasoning. The detailed prompts designed for the interactive graph retriever are provided in Appendix B. An example of interactive reasoning process of Graph-$S^{3}$ can be found in Appendix D.
4 Experiments
Retriever + Generator WebQSP CWQ MetaQA 1-hop MetaQA 2-hop MetaQA 3-hop
Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$
No graph + Qwen3-8B 5.16 8.11 6.26 7.35 2.00 2.88 0.07 0.95 0.20 1.19
No retriever + Qwen3-8B 0.25 1.83 0.37 1.29 0.00 0.29 0.00 0.49 0.00 1.25
RAG/1hop + Qwen3-8B 27.89 38.57 12.55 16.18 75.93 86.36 0.77 1.74 4.13 11.99
RAG/2hop + Qwen3-8B 14.07 24.47 7.00 10.55 42.03 55.59 10.07 21.65 2.60 9.42
RAG/3hop + Qwen3-8B 1.54 7.94 0.99 3.12 0.37 3.03 0.13 2.21 0.13 3.45
ToG + Qwen3-8B 6.14 9.79 7.01 9.61 1.37 1.75 0.00 0.00 0.00 0.20
LightRAG + Qwen3-8B 18.39 31.67 16.20 23.09 1.13 1.76 0.00 0.19 0.07 0.40
G-retriever + Qwen3-8B 25.74 35.45 15.38 18.62 0.63 1.60 0.10 0.77 0.03 1.87
Graph-$S^{3}$ + Qwen3-8B 36.24 47.88 17.87 23.29 81.50 90.22 53.73 65.60 12.73 29.49
No graph + Llama3.1-8B 8.97 15.69 9.40 11.58 12.20 17.60 1.27 7.77 1.23 8.86
No retriever + Llama3.1-8B 0.18 1.97 0.14 1.29 0.00 0.58 0.00 1.11 0.00 2.94
RAG/1hop + Llama3.1-8B 24.82 35.28 13.85 17.26 60.17 70.84 2.40 5.98 4.03 15.46
RAG/2hop + Llama3.1-8B 11.06 22.94 6.29 10.68 29.07 42.47 4.50 15.06 1.80 11.30
RAG/3hop + Llama3.1-8B 1.04 6.67 0.65 3.43 0.33 3.67 0.17 3.37 0.07 5.76
ToG + Llama3.1-8B 8.85 14.28 8.42 12.33 12.40 15.88 0.00 0.63 1.43 6.10
LightRAG + Llama3.1-8B 15.85 36.66 8.33 15.01 13.13 21.47 0.93 4.33 1.00 6.38
G-retriever + Llama3.1-8B 22.67 32.26 13.91 17.47 0.67 1.56 0.10 0.83 0.10 1.77
Graph-$S^{3}$ + Llama3.1-8B 32.31 43.26 17.11 21.17 67.50 76.56 40.17 55.49 10.73 29.55
No graph + Finetuned-8B 9.21 14.88 10.17 12.31 1.63 2.49 0.43 2.64 0.33 4.60
No retriever + Finetuned-8B 0.37 2.26 0.68 1.76 0.00 0.29 0.00 0.34 0.00 1.05
RAG/1hop + Finetuned-8B 28.87 41.48 19.85 26.01 59.50 69.54 1.83 7.44 3.30 18.61
RAG/2hop + Finetuned-8B 14.93 27.76 8.89 14.56 35.83 52.03 7.70 21.21 3.07 14.04
RAG/3hop + Finetuned-8B 1.54 7.81 0.93 4.36 0.57 2.99 0.43 2.31 0.13 4.23
ToG + Finetuned-8B 5.04 9.43 7.54 9.79 2.23 3.44 0.00 0.12 0.10 2.80
LightRAG + Finetuned-8B 17.38 32.59 13.85 19.96 13.77 20.60 0.97 3.71 0.53 4.28
G-retriever + Finetuned-8B 30.34 43.49 22.68 28.38 8.93 11.51 2.33 4.80 0.40 3.31
Graph-$S^{3}$ + Finetuned-8B 44.29 58.45 23.62 30.73 82.77 92.04 63.17 76.18 14.70 36.29
Table 1: Overall results on graph-based QA benchmarks. The best results are highlighted in bold and the second performance results are indicated by an underscore.
4.1 Experimental Setup
Datasets. We evaluate Graph-$S^{3}$ on three widely used textual graph QA benchmarks. WebQSP (Yih et al., 2015) consists of real-world questions annotated with SPARQL queries against Freebase, primarily involving one- or two-hop reasoning. CWQ (Talmor and Berant, 2018) extends WebQSP with more complex multi-hop questions, posing a greater challenge for long reasoning chains. MetaQA (Zhang et al., 2018) is a movie-domain benchmark containing 135k triples and 43k entities, designed to evaluate multi-hop reasoning in a closed domain. Following prior work (Chen et al., 2024), we report accuracy (Acc) and $\mathbf{F_{1}}$ score as evaluation metrics.
Baselines. To validate the effectiveness of our approach, we compare with several representative graph retrieval methods. We additionally evaluate the model’s inherent graph understanding capability through two configurations: (1) the no graph setting, where the model processes the query without any graph input, and (2) the no retriever setting, where the model receives the entire graph structure directly as input.
For traditional RAG, we implemented a multi-hop method where the model retrieves the most relevant graph nodes for the current query and then performs a k-hop expansion to collect information for answer generation. We further compare with representative graph retrievers, including Think-on-Graph (ToG) (Sun et al., 2023), LightRAG (Guo et al., 2024b), G-Retriever (He et al., 2024).
Implementation Details are provided in the Appendix A.
4.2 Main Results
The results of WebQSP, CWQ, and MetaQA are summarized in Table 1. In general, our framework achieves the best performance among all compared methods, demonstrating the effectiveness of combining two-stage training with interactive retrieval1. We break down the analysis into several key observations.
Superiority in Complex Multi-hop Reasoning. Graph-$S^{3}$ exhibits a widening performance gap over baselines as reasoning complexity increases. On the MetaQA benchmark, while competitive on 1-hop questions, our method dominates on multi-hop tasks. Specifically, on MetaQA 3-hop, Graph-$S^{3}$ significantly outperforming the next best agentic baseline and far surpassing traditional $k$-hop RAG methods which struggle to maintain coherence over long chains.
Effectiveness Against Training-Free Baselines. Against sophisticated training-free baselines like ToG and LightRAG, our model delivers substantial gains across all datasets. This highlights that while LLM-based planning (ToG) or structural awareness (LightRAG) is beneficial, it is insufficient without the dedicated fine-tuning and reinforcement learning provided by our pipeline to align the retriever with graph-specific reasoning patterns.
Advantage Over Trained Retrievers. Compared to supervised retrievers like G-Retriever, Graph-$S^{3}$ consistently achieves higher $F_{1}$ scores, indicating better precision in answer generation. This improvement suggests that our interactive pruning mechanism effectively balances recall and precision—retrieving enough context to answer correctly without overwhelming the generator with noise, a common pitfall for G-Retriever which retrieves larger, noisier subgraphs.
Robustness Across Backbones. The performance gains are consistent across different LLM backbones and fine-tuning strategies. Notably, even when using a fine-tuned backbone, Graph-$S^{3}$ maintains its lead, achieving the highest scores on all benchmarks. This demonstrates that our method’s benefits are orthogonal to the underlying model’s capability and can serve as a universal enhancement for graph-based QA systems.
To gain deeper insights into the quantitative improvements, we analyzed specific reasoning examples to verify the effectiveness of our interactive pruning and stepwise supervision. A detailed presentation of these case studies and the corresponding reasoning processes of Graph-$S^{3}$ can be found in Appendix C.
4.3 In-depth Analysis
Methods of Ablation Dataset
WebQSP CWQ MetaQA
1hop 2hop 3hop
Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$
Graph-$S^{3}$ 44.29 58.45 23.62 30.73 82.77 92.04 63.17 76.18 14.70 36.29
w/o SFT 31.64 44.41 7.74 8.77 81.27 89.38 46.30 54.12 2.07 4.98
w/o RL 41.77 53.02 13.39 15.97 71.97 80.09 35.93 45.25 5.73 11.46
w/o interactive 28.87 41.48 19.85 26.01 59.50 69.54 1.83 7.44 3.30 18.61
w/o traj. refine 16.46 19.24 4.12 4.87 39.47 41.06 4.01 6.10 1.34 1.80
Table 2: Results of ablation studies.
4.3.1 Ablation Study
Our framework consists of four key components: supervised fine-tuning (SFT), reinforcement learning (RL) with stepwise rewards, interactive retrieval at inference time, and trajectory refinement during data synthesis. To assess the contribution of each component, we remove one module at a time and evaluate the resulting performance degradation. The results are reported in Table 2.
Ablation of SFT. Removing the SFT stage leads to a clear drop in Accuracy and $\mathbf{F_{1}}$ across all benchmarks. This confirms that SFT provides the retriever with essential navigation ability, compensating for the lack of graph-specific training during upstream pretraining and establishing a stable foundation for subsequent RL optimization.
Ablation of RL. Eliminating the RL stage results in consistent performance degradation, with particularly large declines on CWQ and MetaQA, which require longer reasoning chains. This demonstrates that RL with stepwise rewards substantially strengthens the retriever’s reasoning capability, especially on complex multi-hop tasks.
Ablation of interactive inference. Disabling interactive retrieval causes significant performance drops on 2-hop and 3-hop questions, where results approach those of conventional $k$-hop RAG. This shows that interactive retrieval is crucial for adaptively controlling retrieval depth, effectively filtering redundant neighbors while preserving critical relations.
Ablation of trajectory refinement. Removing trajectory refinement during data synthesis leads to the largest degradation among all ablations. The results indicate that without refinement, synthetic trajectories contain redundant detours, which produce noisy reward signals and undermine the stability of RL optimization.
Retriever Train Method Dataset
WebQSP CWQ MetaQA
1hop 2hop 3hop
Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$ Acc $\mathbf{F_{1}}$
w/o step supervision 41.83 53.87 13.47 16.40 72.63 81.12 34.97 45.14 6.43 11.34
Graph-$S^{3}$ 44.29 58.45 23.62 30.73 82.77 92.04 63.17 76.18 14.70 36.29
Table 3: Performance comparison of process-level rewards and outcome-based rewards training methods.
4.3.2 Effectiveness of Stepwise Supervision
To further validate the effectiveness of our proposed stepwise supervision, we conducted an ablation study. Specifically, starting from the SFT-trained model, we ablated the stepwise reward signals and modified the setup to rely solely on outcome-based rewards. The results of this ablation study are shown in Table 3. Experimental results indicate that without stepwise rewards, model performance experiences a significant decline across all benchmarks, particularly on CWQ and MetaQA which involve longer reasoning chains. This confirms that fine-grained stepwise supervision enables more stable optimization and better generalization on complex multi-hop reasoning tasks.
Figure 3: Number of retrieved graph triples in Graph-$S^{3}$ and baselines on correct answers.
4.3.3 Effective Information Quantification Analysis
To evaluate the efficiency of Graph-$S^{3}$ in retrieving concise yet effective information, we compare it with baseline approaches by measuring the number of triples required to produce correct answers (see Figure 3). Unlike traditional methods that often retrieve large amounts of redundant information, our approach significantly reduces retrieval size while maintaining higher accuracy. In particular, our experiments show that Graph-$S^{3}$ requires only 11.44% of the triples retrieved by G-Retriever on average, yet still achieves superior accuracy. These results highlight the framework’s ability to balance search depth with precision, thereby reducing redundancy.
5 Conclusion
We investigated the limitations of existing retrieval-augmented generation methods on textual graphs, highlighting their reliance on outcome-based supervision and their tendency to produce redundant or incomplete subgraphs. To overcome these challenges, we proposed a framework that integrates three key innovations:
(1) A pipeline for high-quality, stepwise-supervised data synthesis; (2) Two-stage training (SFT then RL) with process-level rewards; (3) Fine-grained, interactive retrieval over textual graphs. Extensive experiments on WebQSP, CWQ, and MetaQA demonstrate that our approach consistently improves both accuracy and $F_{1}$, validating the effectiveness of synthetic stepwise supervision and the proposed training strategy for enhancing interactive graph retrieval. These findings underscore the potential of fine-grained process supervision in bridging the gap between LLMs and structured knowledge, paving the way for more reliable and interpretable agentic reasoning systems.
6 Limitation
While this work has focused on validating Graph-S³ on standard textual graph benchmarks, future research could explore extending the interactive retrieval framework to diverse problem scenarios. Additionally, with the rapid advancement of agent technology, the current framework can be further enhanced by designing a broader spectrum of actions and integrating external tools to tackle increasingly complex reasoning tasks.
7 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62272261, Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120, Tsinghua University (AIR)–AsiaInfo Technologies (China) Inc. Joint Research Center and a grant from GDS Holdings Limited.
Appendix A Training Details
Implementation Details. In our experiments, we primarily employed the Llama3.1-8B (Dubey et al., 2024) and Qwen3-8B (Yang et al., 2025) models. Our data synthesis pipeline produced 9,035 SFT and 3,504 RL training instances; with this data, the Qwen3-8B model was trained for 3 SFT and 15 RL epochs on 8 A100 GPUs, requiring 32 hours in total.
For data generation, we apply our proposed data synthesis pipeline, producing a total of 9,035 training instances for SFT and 3,504 instances for RL. In the SFT stage, we fine-tune the Qwen3-8B with a learning rate of $1\times 10^{-4}$ for $3$ epochs. In the RL stage, we adopt the GRPO algorithm with a batch size of $512$, $15$ training epochs, a learning rate of $1\times 10^{-5}$, a value clipping range of $0.5$, and a KL divergence coefficient of $0.001$. The entire RL training phase takes approximately $32$ hours on $8$ NVIDIA A100 80GB GPUs.
Hyperparameter Value
Learning rate $1\times 10^{-5}$
Batch size 512
Epochs 15
Clip ratio 0.2
Gradient clipping 1.0
KL coefficient 0.001
PPO mini-batch size 16
Table 4: Key hyperparameters for RL training (GRPO).
Appendix B Prompt of Interactive Graph Retrieval
Prompts for Interactive Graph Retriever You are an intelligent agent skilled in exploring Knowledge Graphs, with strong reasoning abilities. Your task is to perform question answering over a Knowledge Graph by gradually exploring it. You should start from the entities mentioned in the question and explore the graph step by step until you gather enough information to answer the question. Your task follows these steps: 1. Understand the Question 2. Analyze the Action History and Current Graph State 3. Choose the Next Action from the following options: "Explore Entity": Explore all triples directly connected to a given entity in the Knowledge Graph. "Choose Relation": Select the triple(s) from the explored information that are most relevant to the question.
Attention: Only the triples included in the "Objects" field of the "Choose Relation" step will be retained in the future "Current Graph State". So You must filter and retain the information useful for answering the question or for further exploration.
"Finish": Choose this action when you believe you have gathered sufficient information to answer the question. Your final answers should be included in the "Objects" field.
4. Select the Objects: Depending on the action, provide the relevant entity or triple(s).
Attention: All objects must come from the "Entities in Question" or the current "Current Graph State". Do not create new entities or relations.
Examples: If "Explore Entity": "Objects": ["EntityA", "EntityB"] If "Choose Relation": "Objects": ["(Sub1, Rel1, Obj1)", ...] If "Finish": "Objects": ["Answer1", "Answer2"]
5. Output your response in JSON format, and include a detailed thought process explaining your reasoning at this step.
  Question:
Entities in Question:
Current Graph State:
Action History:
  Please respond using the following format: Thought Process:
<Provide a step-by-step analysis>
Action Decision:
‘‘‘json
"Action": "<The type of action: ’Explore Entity’ | ’Choose Relation’ | ’Finish’>",
"Objects": [<The entities or triples>]
Appendix C Case Study
Case 1 Question: When were the movies directed by [Sarah Polley] released? True Answer: 2011, 2006, 2012 Action History: Step 1, Explore Entity, Objects: [’Sarah Polley’] Step 2, Choose Relation, Objects: [’(Away from Her, directed_by, Sarah Polley)’, ’(Stories We Tell, directed_by, Sarah Polley)’, ’(Take This Waltz, directed_by, Sarah Polley)’] Step 3, Explore Entity, Objects: [’Stories We Tell’, ’Away from Her’, ’Take This Waltz’] Step 4, Choose Relation, Objects: [’(Away from Her, release_year, 2006)’, ’(Stories We Tell, release_year, 2012)’] Step 5, Explore Entity, Objects: [’Take This Waltz’] Step 6, Choose Relation, Objects: [’(Take This Waltz, release_year, 2011)’] Step 7, Finish, Answer: [’2006’, ’2011’, ’2012’] Retrieved Triples: (Away from Her, release_year, 2006), (Away from Her, directed_by, Sarah Polley), (Take This Waltz, release_year, 2011), (Stories We Tell, directed_by, Sarah Polley), (Stories We Tell, release_year, 2012), (Take This Waltz, directed_by, Sarah Polley)
To demonstrate the efficacy of our interactive pruning mechanism against high-degree entities, we analyze the query: "When were the movies directed by [Sarah Polley] released?" The entity "Sarah Polley" presents a significant challenge as a "hub node" with dual roles: she is a prolific actress with numerous connections but has directed comparatively fewer films.
Graph-$S^{3}$ Success via Structure-Aware Pruning: As shown in the trajectory, Graph-$S^{3}$ effectively overcomes the noise of Sarah Polley’s acting career. In Step 2, the agent leverages the "Choose Relation" action to strictly select triples containing the "directed_by" relation, while deliberately pruning the dense subgraph of movies where she appears merely as an actress. This aligns with our framework’s design to "avoid introducing redundant context into the LLM" and demonstrates the value of stepwise supervision in learning to distinguish task-relevant paths from high-degree noise. Consequently, the agent precisely navigates to the correct release years.
In contrast, baseline methods exhibited failures characteristic of shallow embedding similarity. The standard RAG suffered from severe lexical confusion. It retrieved unrelated entities such as "Pollyanna" and "Polly of the Circus", erroneously matching the token "Polly" with "Polley." This confirms that similarity-based matching often yields noisy results by ignoring precise relational constraints. G-Retriever failed to filter the entity’s large neighborhood, retrieving generic attributes instead of the specific directorial path. This illustrates the limitation of one-shot retrievers that retrieve large neighborhoods discarding the graph’s relational structure, whereas Graph-$S^{3}$’s interactive policy successfully identified the sparse signal within a noisy subgraph.
Case 2 Question: The screenwriter [Jane Anderson] co-wrote movies with whom? True Answer: Anne Heche, Terry Ryan Action History: Step 1, Explore Entity, Objects: [’Jane Anderson’] Step 2, Choose Relation, Objects: [’(If These Walls Could Talk 2, written_by, Jane Anderson)’, ’(Normal, written_by, Jane Anderson)’, ’(The Prize Winner of Defiance, Ohio, written_by, Jane Anderson)’] Step 3, Explore Entity, Objects: [’If These Walls Could Talk 2’, ’Normal’, ’The Prize Winner of Defiance, Ohio’] Step 4, Choose Relation, Objects: [’(If These Walls Could Talk 2, written_by, Jane Anderson)’, ’(Normal, written_by, Jane Anderson)’, ’(The Prize Winner of Defiance, Ohio, written_by, Jane Anderson)’] Step 5, Finish, Answer: [’If These Walls Could Talk 2’, ’Normal’, ’The Prize Winner of Defiance, Ohio’] Retrieved Triples: (The Prize Winner of Defiance, Ohio, written_by, Jane Anderson), (Normal, written_by, Jane Anderson), (If These Walls Could Talk 2, written_by, Jane Anderson)
To identify limitations in current reasoning policies, we analyze a failure case involving the collaborative query: "The screenwriter [Jane Anderson] co-wrote movies with whom?" The trajectory of the agent reveals a critical flaw in the processing of "sibling" relationships. In Steps 1 through 3, the agent correctly identified the intermediate nodes, successfully retrieving movies written by Jane Anderson.
The reasoning collapsed at Step 4. When expanding from the movie nodes to identify other writers, the agent exhibited a strong "self-verification bias." Instead of selecting edges pointing to new entities (the co-writers, such as Anne Heche), the agent selected the "written_by" relations pointing back to the query entity, Jane Anderson. Consequently, instead of discovering the target "sibling" nodes, the agent entered a redundant verification loop. It merely reconfirmed that Jane Anderson wrote the movies and prematurely terminated the search, outputting the movie titles instead of the co-writers’ names. The model assigns disproportionate attention weight to entities explicitly named in the query, causing it to overlook unmentioned but structurally relevant entities.
Appendix D An Example of The Reasoning Process of Graph-$S^{3}$
"step": 0,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"now_state": [],
"action_history": [],
"extract_res": {
"Action": "Explore Entity",
"Objects": [
"The Life of Oharu"
"step": 1,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"now_state": [
"(The Life of Oharu, starred_actors, Kinuyo Tanaka)",
"(The Life of Oharu, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, release_year, 1952)",
"(The Life of Oharu, written_by, Kenji Mizoguchi)",
"(The Life of Oharu, has_tags, kenji mizoguchi)",
"(The Life of Oharu, in_language, Japanese)"
"action_history": [
"step 1, Explore Entity, Objects: [’The Life of Oharu’]"
"extract_res": {
"Action": "Choose Relation",
"Objects": [
"(The Life of Oharu, directed_by, Kenji Mizoguchi)"
"step": 2,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"now_state": [
"(The Life of Oharu, directed_by, Kenji Mizoguchi)"
"action_history": [
"step 1, Explore Entity, Objects: [’The Life of Oharu’]",
"step 2, Choose Relation, Objects: [’(The Life of Oharu, directed_by, Kenji Mizoguchi)’]"
"extract_res": {
"Action": "Explore Entity",
"Objects": [
"Kenji Mizoguchi"
"step": 3,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"now_state": [
"(Ugetsu, directed_by, Kenji Mizoguchi)",
"(Sansho the Bailiff, directed_by, Kenji Mizoguchi)",
"(The Crucified Lovers, directed_by, Kenji Mizoguchi)",
"(The 47 Ronin, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, directed_by, Kenji Mizoguchi)",
"(Osaka Elegy, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, written_by, Kenji Mizoguchi)",
"(Sisters of the Gion, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, has_tags, kenji mizoguchi)",
"(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)"
"action_history": [
"step 1, Explore Entity, Objects: [’The Life of Oharu’]",
"step 2, Choose Relation, Objects: [’(The Life of Oharu, directed_by, Kenji Mizoguchi)’]",
"step 3, Explore Entity, Objects: [’Kenji Mizoguchi’]"
"extract_res": {
"Action": "Choose Relation",
"Objects": [
"(Ugetsu, directed_by, Kenji Mizoguchi)",
"(Sansho the Bailiff, directed_by, Kenji Mizoguchi)",
"(The Crucified Lovers, directed_by, Kenji Mizoguchi)",
"(The 47 Ronin, directed_by, Kenji Mizoguchi)",
"(Osaka Elegy, directed_by, Kenji Mizoguchi)",
"(Sisters of the Gion, directed_by, Kenji Mizoguchi)",
"(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)"
"step": 4,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"now_state": [
"(Ugetsu, directed_by, Kenji Mizoguchi)",
"(Sansho the Bailiff, directed_by, Kenji Mizoguchi)",
"(The Crucified Lovers, directed_by, Kenji Mizoguchi)",
"(The 47 Ronin, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, directed_by, Kenji Mizoguchi)",
"(Osaka Elegy, directed_by, Kenji Mizoguchi)",
"(Sisters of the Gion, directed_by, Kenji Mizoguchi)",
"(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)"
"action_history": [
"step 1, Explore Entity, Objects: [’The Life of Oharu’]",
"step 2, Choose Relation, Objects: [’(The Life of Oharu, directed_by, Kenji Mizoguchi)’]",
"step 3, Explore Entity, Objects: [’Kenji Mizoguchi’]",
"step 4, Choose Relation, Objects: [’(Ugetsu, directed_by, Kenji Mizoguchi)’, ’(Sansho the Bailiff, directed_by, Kenji Mizoguchi)’, ’(The Crucified Lovers, directed_by, Kenji Mizoguchi)’, ’(The 47 Ronin, directed_by, Kenji Mizoguchi)’, ’(Osaka Elegy, directed_by, Kenji Mizoguchi)’, ’(Sisters of the Gion, directed_by, Kenji Mizoguchi)’, ’(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)’]"
"extract_res": {
"Action": "Explore Entity",
"Objects": [
"Ugetsu",
"Sansho the Bailiff",
"The Crucified Lovers",
"The 47 Ronin",
"Osaka Elegy",
"Sisters of the Gion",
"Utamaro and His Five Women"
"step": 5,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"true_answer": [
"Drama"
"now_state": [
"(Ugetsu, directed_by, Kenji Mizoguchi)",
"(Sansho the Bailiff, directed_by, Kenji Mizoguchi)",
"(The Crucified Lovers, directed_by, Kenji Mizoguchi)",
"(The 47 Ronin, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, directed_by, Kenji Mizoguchi)",
"(Osaka Elegy, directed_by, Kenji Mizoguchi)",
"(Sisters of the Gion, directed_by, Kenji Mizoguchi)",
"(Sisters of the Gion, has_genre, Drama)",
"(Ugetsu, has_genre, Drama)",
"(Ugetsu, starred_actors, Masayuki Mori)",
"(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)"
"action_history": [
"step 1, Explore Entity, Objects: [’The Life of Oharu’]",
"step 2, Choose Relation, Objects: [’(The Life of Oharu, directed_by, Kenji Mizoguchi)’]",
"step 3, Explore Entity, Objects: [’Kenji Mizoguchi’]",
"step 4, Choose Relation, Objects: [’(Ugetsu, directed_by, Kenji Mizoguchi)’, ’(Sansho the Bailiff, directed_by, Kenji Mizoguchi)’, ’(The Crucified Lovers, directed_by, Kenji Mizoguchi)’, ’(The 47 Ronin, directed_by, Kenji Mizoguchi)’, ’(Osaka Elegy, directed_by, Kenji Mizoguchi)’, ’(Sisters of the Gion, directed_by, Kenji Mizoguchi)’, ’(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)’]",
"step 5, Explore Entity, Objects: [’Ugetsu’, ’Sansho the Bailiff’, ’The Crucified Lovers’, ’The 47 Ronin’, ’Osaka Elegy’, ’Sisters of the Gion’, ’Utamaro and His Five Women’]"
"extract_res": {
"Action": "Choose Relation",
"Objects": [
"(Ugetsu, has_genre, Drama)",
"(Sisters of the Gion, has_genre, Drama)"
"step": 13,
"question": "the films that share directors with the film [The Life of Oharu] were in which genres",
"question_entities": [
"The Life of Oharu"
"true_answer": [
"Drama"
"now_state": [
"(Ugetsu, directed_by, Kenji Mizoguchi)",
"(Sansho the Bailiff, directed_by, Kenji Mizoguchi)",
"(The Crucified Lovers, directed_by, Kenji Mizoguchi)",
"(The 47 Ronin, directed_by, Kenji Mizoguchi)",
"(The Life of Oharu, directed_by, Kenji Mizoguchi)",
"(Osaka Elegy, directed_by, Kenji Mizoguchi)",
"(Sisters of the Gion, directed_by, Kenji Mizoguchi)",
"(Sisters of the Gion, has_genre, Drama)",
"(Ugetsu, has_genre, Drama)",
"(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)"
"action_history": [
"step 1, Explore Entity, Objects: [’The Life of Oharu’]",
"step 2, Choose Relation, Objects: [’(The Life of Oharu, directed_by, Kenji Mizoguchi)’]",
"step 3, Explore Entity, Objects: [’Kenji Mizoguchi’]",
"step 4, Choose Relation, Objects: [’(Ugetsu, directed_by, Kenji Mizoguchi)’, ’(Sansho the Bailiff, directed_by, Kenji Mizoguchi)’, ’(The Crucified Lovers, directed_by, Kenji Mizoguchi)’, ’(The 47 Ronin, directed_by, Kenji Mizoguchi)’, ’(Osaka Elegy, directed_by, Kenji Mizoguchi)’, ’(Sisters of the Gion, directed_by, Kenji Mizoguchi)’, ’(Utamaro and His Five Women, directed_by, Kenji Mizoguchi)’]",
"step 5, Explore Entity, Objects: [’Ugetsu’, ’Sansho the Bailiff’, ’The Crucified Lovers’, ’The 47 Ronin’, ’Osaka Elegy’, ’Sisters of the Gion’, ’Utamaro and His Five Women’]",
"step 6, Choose Relation, Objects: [’(Ugetsu, has_genre, Drama)’, ’(Sisters of the Gion, has_genre, Drama)’]"
"extract_res": {
"Action": "Finish",
"Objects": [
"Drama"